OpenStudy (anonymous):

What is the root of (x^2+2)?

OpenStudy (anonymous):

break it down. square root of x^2 and then sqrt of 2

OpenStudy (anonymous):

\[ x^2+2=0 \implies x^2=-2 \implies |x| = \sqrt{-2} \]No real roots.